# Unsupervised Detection of Anomalous Driving Patterns Using High Resolution Telematics Time Series Data

### v1

@@ 6.3 Pooled Model
Using the fitted pooled model, we determine the anomaly indices of all trips from all rentals over their entire rental periods, rather than just a 3-day window. This allows us to examine any differences in driving behaviour, focusing particularly on differences in driving anomalies, between rentals with and without (at-fault) claims. The analysis is repeated with four different random training sets, each with a 10% claim rate, and results are reported in 10.
@@
Next, we investigate whether anomaly indices differ when aggregated over entire rental periods for the claimed and no-claim groups. This analysis examines the empirical distribution of anomaly indices for each rental. In the sub-table titled ‘rental-basis’, we present the averages of two metrics for each group: the rental-wise average indices and the rental-wise maximum indices. Specifically, we calculate the average of the rental averages and the average of the rental maximums for each group. While the average indices between the groups remain similar, the claimed group exhibits significantly higher maximums. On average, the maximum indices for the claimed group are roughly double those of the no-claim group across all response dimensions. To help visualize these differences, Figure 3(f) presents the empirical density plots of the anomaly indices for the claimed and no-claim groups on a rental basis. The left column displays the rental averages, while the right column illustrates the rental maximums, with the claimed group represented in red and the no-claim group in blue.
@@ 7 Conclusion
A further direction is to extend the current static/offline training approach by incorporating online learning for HMMs. In this work, models were trained on trip data over relatively short periods of time, such as a 3-day window for individual-specific models in the real-data analysis. Future research can allow the model to dynamically adapt to evolving individual and population driving behaviour. This will enable continuous risk assessment over longer periods and provide more comprehensive insights into long-term driving risks.
@@ Appendix B Controlled Study: Anomaly Indices from Individual-Specific Models with Fewer States
Table 13: UAH-DriveSet: Anomaly index in each dimension of telematics observations, computed from individual-specific CTHMMs with 5 states for each driver and averaged over 100 random initializations. Note: S stands for Secondary Road and M stands for Motorway; Z and Y are longitudinal and lateral accelerations, respectively.
@@
Table 14: UAH-DriveSet: Anomaly index in each dimension of telematics observations, computed from individual-specific CTHMMs with 10 states for each driver and averaged over 100 random initializations. Note: S stands for Secondary Road and M stands for Motorway; Z and Y are longitudinal and lateral accelerations, respectively.
@@
Table 15: UAH-DriveSet: Anomaly index in each dimension of telematics observations, computed from individual-specific CTHMMs with 15 states for each driver and averaged over 100 random initializations. Note: S stands for Secondary Road and M stands for Motorway; Z and Y are longitudinal and lateral accelerations, respectively.
@@ Appendix C Real-Data Analysis: Summaries of Logistic Models from Individual-Specific Models
Table 16: Model summary of logistic GLMs, with raw or normalized anomaly indices as covariates. Note: x1, x2 and x3 denote indices for speed, longitudinal and lateral accelerations respectively.
@@
Table 17: Feature importance of logistic XGBoost, with raw or normalized anomaly as covariates. Note: x1, x2 and x3 denote indices for speed, longitudinal and lateral accelerations respectively.
@@ Appendix D Real-Data Analysis: Pooled Model Analysis with Alternative Training Sets
Table 18: Comparison of anomaly indices (%) in each telematics response dimension for claimed and no-claim rentals, computed from pooled CTHMMs trained on samples with 0% and 100% claim rates.
@@
Table 19: 5-fold cross-validation ROC-AUC of logistic GLMs with different sets of covariates. Set 1: maximum anomaly indices, Set 2: right tail of the anomaly index empirical distribution, Set 3: right tail and the number of trips as an offset. Note: first column indicates the claim rate of the training set the pooled CTHMM is fitted on.

### v2

@@ 6.3 Pooled Model
Using the fitted pooled model, we determine the anomaly indices of all trips from all rentals over their entire rental periods, rather than just a 3-day window. This allows us to examine any differences in driving behaviour, focusing particularly on differences in driving anomalies, between rentals with and without (at-fault) claims. The analysis is repeated with four different random training sets, each with a 10% claim rate, and results are reported in Table 13.
@@
Next, we investigate whether anomaly indices differ when aggregated over entire rental periods for the claimed and no-claim groups. This analysis examines the empirical distribution of anomaly indices for each rental. In the sub-table titled ‘rental-basis’, we present the averages of two metrics for each group: the rental-wise average indices and the rental-wise maximum indices. Specifically, we calculate the average of the rental averages and the average of the rental maximums for each group. While the average indices between the groups remain similar, the claimed group exhibits significantly higher maximums. On average, the maximum indices for the claimed group are roughly double those of the no-claim group across all response dimensions. To help visualize these differences, Figure 4 presents the empirical density plots of the anomaly indices for the claimed and no-claim groups on a rental basis. The left column displays the rental averages, while the right column illustrates the rental maximums, with the claimed group represented in red and the no-claim group in blue.
@@ 7 Conclusion
A further direction is to extend the current static/offline training approach by incorporating online learning for HMMs. In this work, models were trained on trip data over relatively short periods of time, such as a 3-day window for individual-specific models in the real data analysis. Future research can allow the model to dynamically adapt to evolving individual and population driving behaviour. This will enable continuous risk assessment over longer periods and provide more comprehensive insights into long-term driving risks.
@@
Acknowledgement
The authors acknowledge the financial support provided by the Natural Sciences and Engineering Research Council of Canada [RGPIN 284246, RGPIN-2023-04326] and Canadian Institute of Actuaries [CS000271].
Conflicts of interest or Competing interests
The authors declare no conflicts of interest or competing interests in this paper, with no financial or personal affiliations that could compromise the objectivity or integrity of the presented work.
Data Availability
The UAH-DriveSet used in the controlled study is publicly available, while the rental car dataset studied in the real data analysis is proprietary data from a telematics company that we are unable to share.
@@ Appendix B Controlled Study: Anomaly Indices from Individual-Specific Models with Fewer States
Table B1: UAH-DriveSet: Anomaly index in each dimension of telematics observations, computed from individual-specific CTHMMs with 5 states for each driver and averaged over 100 random initializations. Note: S stands for Secondary Road and M stands for Motorway; Z and Y are longitudinal and lateral accelerations, respectively.
@@
Table B2: UAH-DriveSet: Anomaly index in each dimension of telematics observations, computed from individual-specific CTHMMs with 10 states for each driver and averaged over 100 random initializations. Note: S stands for Secondary Road and M stands for Motorway; Z and Y are longitudinal and lateral accelerations, respectively.
@@
Table B3: UAH-DriveSet: Anomaly index in each dimension of telematics observations, computed from individual-specific CTHMMs with 15 states for each driver and averaged over 100 random initializations. Note: S stands for Secondary Road and M stands for Motorway; Z and Y are longitudinal and lateral accelerations, respectively.
@@ Appendix C Real-Data Analysis: Summaries of Logistic Models from Individual-Specific Models
Table C1: Model summary of logistic GLMs, with raw or normalized anomaly indices as covariates. Note: x1, x2 and x3 denote indices for speed, longitudinal and lateral accelerations respectively.
@@
Table C2: Feature importance of logistic XGBoost, with raw or normalized anomaly as covariates. Note: x1, x2 and x3 denote indices for speed, longitudinal and lateral accelerations respectively.
@@ Appendix D Real-Data Analysis: Pooled Model Analysis with Alternative Training Sets
Table D1: Comparison of anomaly indices (%) in each telematics response dimension for claimed and no-claim rentals, computed from pooled CTHMMs trained on samples with 0% and 100% claim rates.
@@
Table D2: 5-fold cross-validation ROC-AUC of logistic GLMs with different sets of covariates. Set 1: maximum anomaly indices, Set 2: right tail of the anomaly index empirical distribution, Set 3: right tail and the number of trips as an offset. Note: first column indicates the claim rate of the training set the pooled CTHMM is fitted on.